2.1x+45.2=3.2-8.4x
Answer by checkley77(12844)

2.1x+45.2=3.2-8.4x
2.1X+8.4X=3.2-45.2
10.5X=-42
X=-42/10.5
X=-4 ANSWER.
PROOF:
2.1*-4+45.2=3.2-8.4*-4
-8.4+45.2=3.2+33.6
36.8=36.8